OpenStudy (firejay5):

A contractor must pour a cement slab foundation for a new home with the floor plan shown. What is the area of the foundation? *Show work and explain answer choice* A. 320ft squared B. 400ft squared C. 520ft squared D. 600ft squared

OpenStudy (firejay5):



OpenStudy (firejay5):

@jdoe0001 @timo86m

OpenStudy (anonymous):

(20 * 80) + ((20 - 8) * 10)

OpenStudy (jdoe0001):



OpenStudy (anonymous):

Does my way work? o.o

OpenStudy (jdoe0001):

well... 20 * 80... is ... a bit large, maybe it was a typo

OpenStudy (anonymous):

20*20 = 400 10*12=120 400+120 = 520